Determine the mole fraction of iron(III) nitrate in a 3.99 M aqueous solution of iron(III) nitrate. The density of the solution is 1.05 g mL-1.
1
Expert's answer
2019-11-08T07:48:49-0500

Mol. Wt of Fe(NO3)3=241.86 gm

3.99 M = 241.86*3.99

= 965.02 gm

Suppose 1000 ml solution.

Weight of the sol 1050 gm

Weight of water= 1050-965.02

= 84.98 gm

=84.98/18 mol

=4.72 mol

Mole fraction of Fe(NO3)3 = 3.99/(3.99+4.72)

=0.458